# Pushing Toward the Simplex Vertices: A Simple Remedy for Code Collapse in Smoothed Vector Quantization

### v1

@@ 1 Introduction
Vector quantization is a method for discretizing a continuous vector space (Gray,, 1984; van den Oord et al.,, 2017). It maps each vector in the continuous space to the nearest element of a finite set of representative vectors (a.k.a. a codebook). The resulting discrete representations are easier to manipulate and interpret than the original continuous forms, and have proven effective across diverse applications, including image generation (Esser et al.,, 2021; Ramesh et al.,, 2021; Rombach et al.,, 2022; Yu et al.,, 2022), speech recognition (Baevski et al., 2020a, ; Baevski et al., 2020b, ), and music generation (Hadjeres and Crestel,, 2020; Dhariwal et al.,, 2020).
@@ 2.3 Regularization
It should be noted, however, that the codebook loss in Equation 10 does not inherently prevent code collapse. Some codebook vectors may never serve as the nearest neighbor of any pre-quantized feature and therefore receive no updates. Accordingly, previous studies have resorted to additional workarounds; for example, unused codebook vectors may be reset to the positions of pre-quantized features Dhariwal et al., (2020).
@@ 4 Experiments
The proposed regularization for smoothed vector quantization was benchmarked on two tasks: discrete autoencoding (§4.1) and contrastive learning (§4.2). The Python code used for these experiments will be published when this manuscript is accepted for a journal/conference.
@@ 4.1 Discrete Autoencoding
Extending prior work (Esser et al.,, 2021; Fifty et al.,, 2025), three different combinations of feature map and codebook sizes were examined. In addition to the previously used settings of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), an additional configuration increased the channel dimensionality of the latter to 3232 ($H\times W\times C=64\times 64\times 32,M=8196$).
Each model was trained using four GPUs, and the proposed method computed the $K/4$-nearest neighbors of each simplex vertex per GPU (see §5.3 for further discussion). The number of neighbors was set to $K=2\times 4$ for the spatially smaller feature maps of size $16\times 16\times 32$, and to $K=8\times 4$ for the larger ones. [^3] The value of $K$ was upper-bounded at $8\times 4$ by available computational resources; With a maximum batch size of $64$, the total number of latent pixels was $64\times 64\times 64=8\times 4\times 8196$, allowing only $8\times 4$ neighbors per vertex of $\Delta^{8196}$. This implementation constraint is further discussed in §5.3.
Table 1: Performance of discrete autoencoding on the ImageNet validation set. Reported metrics are codebook usage and reconstruction quality scores: root mean squared error (rMSE), Fréchet Inception Distance (FID), and Inception Score (IS). The proposed method is denoted as “KNN-L2/CE”.
Method Feature Map Size; Codebook Size
@@
Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS
@@
PPL 100.0% 1.112 309.603 4.552 81.6% 0.701 37.987 196.723 99.8% 0.725 83.070 77.981
KNN-L2 100.0% 0.366 81.691 71.750 100.0% 0.199 12.644 361.228 100.0% 0.204 7.908 381.654
KNN-CE 100.0% 0.380 106.048 50.750 100.0% 0.219 14.241 343.228 100.0% 0.199 5.267 404.098
Table 3 reports the codebook usage and reconstruction quality scores—including root mean squared error (rMSE), Inception Score (IS; Salimans et al.,, 2016), and Fréchet Inception Distance (FID; Heusel et al.,, 2017)–for the proposed method (“KNN-L2/CE”) and the baselines. The results indicate that reconstruction through softmax-smoothed quantization (i.e., without Gumbel randomness) deteriorated substantially when combined with the perplexity-based regularization (“PPL”; Equation 11). This degradation stems from the mismatch between soft quantization during the training and hard quantization at inference, supporting the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing (i.e., vertex-concentrated distributions).
The perplexity-based regularization was only effective when combined with the Gumbel-softmax sampling, and onehot quantization in the forward computation by STE was necessary to ensure full codebook usage (“Hard-Gumbel”). Otherwise, code collapse was not prevented when the channel dimensionality was large ($C=64$; “Soft-Gumbel”).
@@
It is also worth noting that both STE and rotational gradient estimation (RE) suffered from code collapse when the channel dimensionality was large ($C=64$), illustrating the difficulty of achieving robust prevention of code collapse across different settings. The next section further highlights this challenge under a different training objective.
@@ 4.2 Contrastive Learning
Two codebook configurations were investigated. The first used a single codebook of size 1024. By contrast, the second followed the original work of Baevski et al., 2020b , combining two smaller codebooks—each of size 320—to implement rich code diversity efficiently via product quantization (Jégou et al.,, 2011, see §5.3 for more information). The dimensionality of the codebook vectors was set to 256 for the single-codebook configuration and to 128 for the dual-codebook configuration. The number of nearest neighbors was set to $K=2\times 4$, meaning that two neighbors were identified per codebook entry on each of the four GPUs. In the dual-codebook condition, the two codebooks were treated independently.
@@
KNN-L2 60.4% 100.0% 100.0%
KNN-CE 100.0% 100.0% 100.0%
Table 4 reports codebook usage. The perplexity-based regularization failed to prevent code collapse in both single- and dual-codebook settings. Likewise, the STE and rotational gradient estimation approaches exhibited the same failure.
By contrast, the proposed KNN-based regularization ensured complete code utilization in both conditions when cross-entropy was used as the divergence metric on the simplex. The L2 metric, on the other hand, proved insufficient to prevent code collapse, especially when combined with Gumbel sampling.
@@ 5.3 Limitations
A primary limitation of the proposed method is its memory requirement. When training across multiple GPUs, the KNN-based regularization identifies $K$ nearest smoothed quantizers ($\mathbf{p}$ or $\boldsymbol{\pi}$) per simplex vertex on each GPU, rather than finding global neighbors across all GPUs. Consequently, each GPU must have sufficient VRAM to store at least $KM$ latent pixels/frames, where $M$ denotes the codebook size. This requirement can become prohibitive when $M$ is large (i.e., for fine-grained quantization).
@@ Acknowledgments
This study was supported by JSPS Grant-in-Aid for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); JST AIP Accelerated Program (JPMJCR25U6) and Core Research for Evolutional Science and Technology (JPMJCR22P5); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620). The author also gratefully acknowledges the support of the Academic Center for Computing and Media Studies, Kyoto University, regarding the use of their supercomputer system.
@@ A.1 Smoothed Quantization
Smoothed quantizers $\mathbf{p}$ were computed as $\mathbf{p}=\operatorname{softmax}(\hat{Q}^{\mathsf{T}}\hat{\mathbf{z}}/\tau)$. In other words, the codebook vectors $(\mathbf{q}_{1},\dots,\mathbf{q}_{M})=Q$ and the feature vectors $\mathbf{z}$ were first L2-normalized, and their product (i.e., cosine similarity) was rescaled by a learnable temperature $\tau$. This temperature was shared across the codebook so that all the logits had the same amplitude. Assignment probabilities $\mathcal{\pi}$ for Gumbel-softmax sampling were computed in the same way. [^4] Previous studies manually annealed the temperature $\tau$ instead of treating it as a learnable parameter (Baevski et al., 2020b, ). This approach was also tested in the experiments here but did not yield improvements over the learnable temperature.
@@ A.2 Hard Quantization
The weight $\beta$ on the commitment loss in Equation 10 was set to 1.0, based on the previous observations that its value does not significantly affect learning outcomes within the range 0.1–2.0. (van den Oord et al.,, 2017).
@@ B.1 Discrete Autoencoding
Both the convolutional encoder and Transformer were implemented using the publicly available code in torchaudio, and only the quantization module was implemented from scratch.

### v2

@@ 1 Introduction
Vector quantization is a method for discretizing a continuous vector space (Gray,, 1984; van den Oord et al.,, 2017). It maps each vector in the continuous space to the nearest element of a finite set of representative vectors (a.k.a. a codebook). The resulting discrete representations are easier to manipulate and interpret than the original continuous forms, and have proven effective across diverse applications, including image generation (Esser et al.,, 2021; Ramesh et al.,, 2021; Rombach et al.,, 2022; Yu et al., 2022b, ), speech recognition (Baevski et al., 2020a, ; Baevski et al., 2020b, ), and music generation (Hadjeres and Crestel,, 2020; Dhariwal et al.,, 2020).
@@ 2.3 Regularization
It should be noted, however, that the codebook loss in Equation 10 does not inherently prevent code collapse. Some codebook vectors may never serve as the nearest neighbor of any pre-quantized feature and therefore receive no updates (Zhu et al.,, 2025). Accordingly, previous studies have resorted to additional workarounds; for example, unused codebook vectors may be reset to the positions of pre-quantized features Dhariwal et al., (2020). More recently, Zhu et al., (2025) proposed another remedy called SimVQ, which reparameterizes the codebook $Q$ as the product of a randomly frozen matrix $Q^{\prime}\in\mathbb{R}^{M\times D}$ and a learnable matrix $W\in\mathbb{R}^{D\times D}$ (i.e., $Q=Q^{\prime}W$). In this formulation, all codebook vectors share learnable parameters with one another ($\mathbf{q}_{m}=q^{\prime}_{m,1}\mathbf{w}_{1}+\cdots+q^{\prime}_{m,D}\mathbf{w}_{D}$), so updates to one vector propagate to the others.
@@ 4 Experiments
The proposed regularization for smoothed vector quantization was benchmarked on two tasks: discrete autoencoding (§4.1) and contrastive learning (§4.2). The Python code used for these experiments is available as supplementary material.
@@ 4.1 Discrete Autoencoding
Extending prior work (Esser et al.,, 2021; Fifty et al.,, 2025), three different combinations of feature map and codebook sizes were examined. In addition to the previously used settings of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), an additional configuration increased the channel dimensionality of the latter to $32$ ($H\times W\times C=64\times 64\times 32,M=8196$).
Each model was trained using four GPUs, and the proposed method computed the $K/4$-nearest neighbors of each simplex vertex per GPU (see §5.3 for further discussion). The number of neighbors was set to $K/4=8$ for the L2 distance and $K/4=1$ for the cross-entropy otherwise specified. [^3] The value of $K$ was upper-bounded at $8\times 4$ by available computational resources; With a maximum batch size of $64$, the total number of latent pixels was $64\times 64\times 64=8\times 4\times 8196$, allowing only $8\times 4$ neighbors per vertex of $\Delta^{8196}$. This implementation constraint is further discussed in §5.3. The weight $\beta$ on the commitment loss (Equation 10) in STE (including SimVQ) and rotational gradient estimation was set to 1.0, following Fifty et al., (2025) and Zhu et al., (2025).
Table 1: Performance of discrete autoencoding on the ImageNet validation set. Reported metrics are codebook usage and reconstruction quality scores: root mean squared error (rMSE), Fréchet Inception Distance (FID), and Inception Score (IS). The proposed method is denoted as “KNN-L2/CE”. Best scores across all methods are highlighted in boldface, while underlined values indicate the best-performing number of nearest neighbors among $K/4\in\{1,2,4,8\}$.
Feature Map Size; Codebook Size
@@
Method $K/4$ Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS
@@
PPL 100.0% 1.112 309.60 4.55 81.6% 0.701 37.99 196.72 99.8% 0.725 83.07 77.98
KNN-L2 1 100.0% 0.358 76.29 81.61 100.0% 0.224 12.08 354.36 100.0% 0.387 32.20 210.61
2 100.0% 0.366 81.69 71.75 100.0% 0.226 12.04 348.48 100.0% 0.175 3.62 427.42
4 100.0% 0.379 87.05 62.37 100.0% 0.205 9.01 383.56 100.0% 0.196 5.64 403.06
8 99.9% 0.343 73.72 88.53 100.0% 0.199 12.64 361.23 100.0% 0.204 7.91 381.65
KNN-CE 1 100.0% 0.366 79.80 74.69 100.0% 0.225 14.08 345.56 100.0% 0.175 2.81 437.72
2 100.0% 0.380 106.05 50.75 100.0% 0.227 14.34 338.37 100.0% 0.187 4.16 418.57
4 100.0% 0.748 255.59 9.03 100.0% 0.228 16.73 326.78 100.0% 0.204 6.98 387.88
8 100.0% 1.240 467.86 3.01 100.0% 0.219 14.24 343.23 100.0% 0.199 5.27 404.10
Table 1 reports the codebook usage and reconstruction quality scores—including root mean squared error (rMSE), Inception Score (IS; Salimans et al.,, 2016), and Fréchet Inception Distance (FID; Heusel et al.,, 2017)–for the proposed method (“KNN-L2/CE”) and the baselines. Reconstruction through softmax-smoothed quantization (i.e., without Gumbel randomness) deteriorated substantially when combined with the perplexity-based regularization (“PPL”; Equation 11). This degradation stems from the mismatch between soft quantization during the training and hard quantization at inference; the smoothed quantizers deviated from the simplex vertices, as reflected in their high individual perplexity, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$ (Table 2). These observations support the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing.
Table 2: Tightness of softmax-based smoothing (without Gumbel sampling), measured by the individual perplexity of smoothed quantizers, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$. Reported values are the 75th, 90th, and 99th percentiles, as well as the maximum, computed across all feature-map pixels in the ImageNet validation set.
Feature Map Size; Codebook Size
$\times16$$\times$32; 1024 $\times64$$\times$3; 8192 $\times64$$\times$32; 8192
Method $K/4$ 75% 90% 99% Max 75% 90% 99% Max 75% 90% 99% Max
PPL 910.40 933.12 963.29 995.91 7559.89 7561.99 7563.10 7565.02 7230.60 7267.06 7315.62 7424.16
KNN-L2 1 1.17 1.59 2.36 6.26 1.00 1.00 1.00 4.00 2.24 40.45 86.53 249.64
2 1.00 1.08 1.86 27.37 1.00 1.00 1.06 4.65 1.14 1.55 2.33 7.92
4 1.00 1.06 1.86 62.99 1.00 1.00 1.00 4.61 1.00 1.06 1.80 20.42
8 1.00 1.10 1.90 120.68 3.33 3.87 5.22 12.03 1.00 1.05 1.79 107.70
KNN-CE 1 1.18 1.60 2.38 6.28 3.96 4.51 5.81 11.96 1.09 1.45 2.17 5.94
2 1.23 1.72 2.73 8.07 4.04 4.66 6.16 11.77 1.06 1.38 2.08 6.15
4 1.42 1.78 2.38 5.45 4.16 4.86 6.57 13.04 1.08 1.44 2.16 6.58
8 2.38 2.80 3.71 7.78 4.46 5.31 7.39 13.81 1.29 1.75 2.65 9.26
The perplexity-based regularization was only effective when combined with the Gumbel-softmax sampling, and onehot quantization in the forward computation by STE was necessary to ensure full codebook usage (“Hard-Gumbel”). Otherwise, code collapse was not prevented when the channel dimensionality was large ($C=32$; “Soft-Gumbel”).
@@
In terms of computational efficiency, however, cross-entropy proved more favorable. Notably, its performance was highest when the number of neighbors was minimal ($K/4=1$) across all feature map sizes (indicated by the underlined scores in Table 1); increasing $K$ did not yield further improvements. Although most smoothed quantizers remained unregularized under this setting, they still achieved low individual perplexities after the training (Table 2), indicating a tight approximation of quantization. By contrast, minimizing only the L2 distance between the simplex vertices and their single nearest neighbor per GPU led to performance degradation for the feature map of size $64\times 64\times 8192$. Since requiring more neighbors necessitates a larger batch size (see §5.3 for details), cross-entropy emerges as a more scalable option.
Finally, both STE and rotational gradient estimation (RE) exhibited severe code collapse when the channel dimensionality was large ($C=32$), reaffirming prior findings that reducing channel dimensionality is necessary for stable training (Yu et al., 2022a, ; Yu et al.,, 2024; Mentzer et al.,, 2024). By contrast, SimVQ maintained full codebook usage across all settings and achieved strong reconstruction quality even at high dimensionality. However, the next section presents a case study in which all gradient-estimation approaches—including SimVQ—encounter code collapse, highlighting the difficulty of achieving robust prevention of code collapse across different settings.
@@ 4.2 Contrastive Learning
Two codebook configurations were investigated. The first used a single codebook of size 1024. By contrast, the second followed the original work of Baevski et al., 2020b , combining two smaller codebooks—each of size 320—to implement rich code diversity efficiently via product quantization (Jégou et al.,, 2011, see §5.3 for more information). The dimensionality of the codebook vectors was set to 256 for the single-codebook configuration and to 128 for the dual-codebook configuration. The weight $\beta$ on the commitment loss in both STE (including SimVQ) and rotational gradient estimation was set to 1.0 (Fifty et al.,, 2025; Zhu et al.,, 2025).
@@
KNN-L2 1 82.4% 100.0% 100.0%
2 60.4% 100.0% 100.0%
KNN-CE 1 99.5% 100.0% 100.0%
2 100.0% 100.0% 100.0%
Table 3 reports codebook usage. The perplexity-based regularization failed to prevent code collapse in both single- and dual-codebook settings. Likewise, the STE and rotational gradient estimation approaches exhibited the same failure. Remarkably, SimVQ—despite achieving full codebook usage in the discrete autoencoding experiments—offered no observable benefit in this learning paradigm.
By contrast, the proposed KNN-based regularization ensured (near-)complete code utilization in both conditions when cross-entropy was used as the divergence metric on the simplex. As in the discrete autoencoding experiments, a single neighbor per GPU was sufficient to obtain this effect. The L2 metric, on the other hand, proved insufficient to prevent code collapse, especially when combined with Gumbel sampling.
@@ 5.3 Limitations
A primary limitation of the proposed method is its memory requirement. When training across multiple GPUs, the KNN-based regularization identifies $K$ nearest smoothed quantizers ($\mathbf{p}$ or $\boldsymbol{\pi}$) per simplex vertex on each GPU, rather than finding global neighbors across all GPUs. Consequently, each GPU must have sufficient VRAM to store at least $KM$ latent pixels/frames, where $M$ denotes the codebook size. This requirement can become prohibitive when $M$ is large (i.e., for fine-grained quantization), although empirically, a single neighbor per GPU appeared sufficient to prevent code collapse when cross-entropy was used as the divergence metric.
@@ Acknowledgments
Hidden for double-blind review.
@@ A.1 Smoothed Quantization
[…]
@@ A.2 Hard Quantization
The weight $\beta$ on the commitment loss in Equation 10 was set to 1.0 (Fifty et al.,, 2025; Zhu et al.,, 2025), based on the previous observations that its value does not significantly affect learning outcomes within the range 0.1–2.0. (van den Oord et al.,, 2017).
@@ B.1 Discrete Autoencoding
Both the convolutional encoder and Transformer were implemented using the publicly available code in torchaudio, and only the quantization module was implemented from scratch. All components are provided in the supplementary material.